# An Extended Closure Relation by LightGBM for Neutrino Radiation Transport in Core-collapse Supernovae

### v1

@@ 4.2 Off-Diagonal Component
Although the diagonal components, particularly the 11-component, is dominant in neutrino transport, the off-diagonal components should not be forgotten (see discussions in Nagakura et al., 2018). Designed to reproduce the dominant component, the M1-closure is known to give large errors to the off-diagonal components sometimes (Harada et al., 2019, 2020). It is hence one of our goals in the machine learning modeling of the Eddington tensor to better reproduce the off-diagonal components. It is admittedly true in this respect that the previous TBNN closure was not so good, either (Harada et al., 2022). In the following, we examine the accuracy of the off-diagonal components in our LightGBM model.
We focus on the 12-component, which is actually the most important off-diagonal component. Figure 20 shows the MAE for this component as a function of the neutrino energy at different postbounce times. We find again that our LightGBM model achieves higher accuracies in general compared to the M1 closure. The radial profiles of the $12$-component of the Eddington tensor are presented in figures 2124. They are obtained for the radial ray at $\theta=\pi/2$. At 250 ms post bounce, our LightGBM model fairs better than the M1 closure except at $\varepsilon_{\nu}$=54.0MeV, where the two models give similar errors. As we go back in time, the deviation from the results of the Boltzmann simulation gets larger for the LightGBM model and its advantage becomes not so remarkable.
Figure 20: The same as figure 4 except that the target variable is $k_{\rm FP}^{12}$.
Figure 21: The same as figure 7 except that the target is $k_{\rm FP}^{12}$.
Figure 22: The same as figure 20 except that the snapshot time is 200 ms.
Figure 23: The same as figure 20 except that the snapshot time is 150 ms.
Figure 24: The same as figure 20 except that the snapshot time is 100 ms.
A potential factor that contributes to the compromized success is the variation in the statistical properties among the data at different time steps. Tables 4 and 5 give the the mean and standard deviation of $k_{\rm boltz,FP}^{11}$ and $k_{\rm boltz,FP}^{12}$ at each time step. It is evident that the statistical properties of $k_{\rm boltz,FP}^{12}$ are significantly different from time to time compared to those of $k_{\rm boltz,FP}^{11}$. It may be that the generalization should not have been expected for the off-diagonal component in the first place. Standardisation of the dataset across all time steps may be a solution but again is out of the scope of this paper.
It is worth noting that different features were utilized for the diagonal and off-diagonal components in our LightGBM model. Specifically for the off-diagonal components, it was unclear which features held the most significance. In predicting the diagonal components, we utilized 35 features, whereas for the off-diagonal components, 317 features were employed. For example, features such as fluid velocity, which were removed when predicting diagonal components, are included. When predicting diagonal components, the prediction accuracy significantly decreased; however, in predicting off-diagonal components, the accuracy was not greatly compromised. Off-diagonal components are more significantly affected by factors such as fluid motion compared to diagonal components, which suggests that background information may be more important for their prediction.
@@ 4.3 Importance of feature engineering
Feature engineering plays a crucial role in predicting the Eddington tensor accurately. In this work we incorporated many advanced features indeed as listed in table 2. It is possible because LightGBM is very efficient both in training and prediction, ignoring irrelevant features automatically. As explained in section 3.2, we find that the spatial-shift and difference features are two most important features for the improvement of accuracy, which we demonstrate here.
In Figure 25, we show how MAE gets better from the one for the model with only the basic features included. The MAE for the model with the basic features alone is actually larger than that for the M1 closure method for most of the neutrino energy. The MAE becomes smaller, however, once the advanced features are incorporated, particularly for energies close to the mean energy. It is further improved if the spatial-shift and difference features are taken into account in addition. The same trend is observed for the Eddington tensor $k_{\rm{FP}}^{11}$ itself as a function of radius as demonstrated in Figure 26.
What we learned from these experiments on feature engineering may be three-fold. Firstly, the local low-order moments, the inputs for the ordinary closure methods, are not sufficient to reproduce the Eddington tensor accurately. We find that the information on the background matter such as the baryon density or the optical depth is particularly important. Secondly, the nonlocal information on the neutrino distribution is probably the most crucial. Although their incorporation in LightGBM is rather straightforward, it will not be so easy for the analytic closure relation to take them into account. Thirdly, finding nice combinations of the basic features is also an important thing. Our LightGBM model is useful in this respect, since we can try many possible combinations in relatively short times. The new features so obtained may be employed in other machine learning models.
Figure 25: The same as the bottom left figure in figure 4 except there are some additional plots: the black one is the LightGBM prediction with basic features (Table 1) and the orange line with advanced features.
Figure 26: LightGBM prediction of $k_{\rm{FP}}^{11}$ with different features $(200\,{\rm ms})$. The blue line corresponds to the Boltzmann-radiation-hydrodynamics code and the green line to the M1 closure method. The black line is the LightGBM prediction with basic features (Table 1), the orange line with advanced features and the red line with all engineered features: advanced features, spatial-shift features and spatial-difference features.
5 Conclusion
@@
Although these successes are encouraging, there are admittedly many unsatisfactory things. For instance, the resultant Eddington factor is not very smooth in radius, which may cause a trouble when we substitute it for the M1 closure relation in the truncated moment method and perform core-collapse simulations. Although it is much improved from our previous neural network model, the generalization performance is not good enough. This may suggest that we should use all data collected from different times for the training. We are currently trying such possibilities. The non-local features investigated in this paper are based on physical intuitions and admittedly very primitive, and there is a much room for improvement. The ultimate goal of this project is to create a versatile model capable of predicting the Eddington tensor (and even higher moments as well) for stars with different masses and/or degrees of rotation. It may turn out that the machine based on the decision tree is not the best one. Even in that case, the features we identify in this work may be used as the inputs to other machine learning models. The current model is useful to find other features to improve the prediction further.
@@ Acknowledgments
We acknowledge Hiroki Nagakura and Keiya Hirashima for fruitful discussions. This work was supported by KAKENHI Grant Numbers JP21K13913, 21H01083. This work was also supported by MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Toward a unified view of the universe: from large scale structures to planets). S. T. is supported by International Graduate Program of Innovation for Intelligent World. S. Y. is supported by Institute for Advanced Theoretical and Experimental Physics, Waseda University and the Waseda University Grant for Special Research Projects (project number: 2023-C141, 2024-C56, 2024-Q014). We acknowledge the high-performance computing resources of the K-computer / the supercomputer Fugaku provided by RIKEN, the FX10 provided by Tokyo University, the FX100 provided by Nagoya University, the Grand Chariot provided by Hokkaido University, and Oakforest-PACS / Wisteria Odyssey provided by JCAHPC through the HPCI System Research Project (Project ID: hp130025, 140211, 150225, 150262, 160071, 160211, 170031, 170230, 170304, 180111, 180179, 180239, 190100, 190160, 200102, 200124, 210050, 220047, 220223, 230056, 230270, 240041) for producing and processing the supervisor data.

### v2

@@ 4.2 Off-Diagonal Component
Although the diagonal components, particularly the 11-component, are dominant in neutrino transport, the off-diagonal components should not be forgotten (see discussions in Nagakura et al., 2018). Designed to reproduce the dominant component, the M1 closure is known to give large errors to the off-diagonal components sometimes (Harada et al., 2019, 2020). It is hence one of our goals in the machine learning modeling of the Eddington tensor to better reproduce the off-diagonal components. It is admittedly true in this respect that the previous TBNN closure was not so good, either (Harada et al., 2022). In the following, we examine the accuracy of the off-diagonal components in our LightGBM model.
We focus on the 12-component, which is actually the most important off-diagonal component. Figure 3 shows the MAE for this component as a function of the neutrino energy at different post-bounce times. We find again that our LightGBM model achieves higher accuracies in general compared to the M1 closure. The radial profiles of the $12$-component of the Eddington tensor are presented in Figure 7. They are obtained for the radial ray at $\theta=\pi/2$. At 250 ms post-bounce, our LightGBM model fares better than the M1 closure except at $\varepsilon_{\nu}$ = 54.0MeV, where the two models give similar errors. As we go back in time, the deviation from the results of the Boltzmann simulation gets larger for the LightGBM model and its advantage becomes not so remarkable.
A potential factor that contributes to the compromised success is the variation in the statistical properties among the data at different time steps. Tables 4 and 5 give the mean and standard deviation of $k_{\rm boltz,FP}^{11}$ and $k_{\rm boltz,FP}^{12}$ at each time step. It is evident that the statistical properties of $k_{\rm boltz,FP}^{12}$ are significantly different from time to time compared to those of $k_{\rm boltz,FP}^{11}$. It may be that the generalization should not have been expected for the off-diagonal component in the first place. Standardisation of the dataset across all time steps may be a solution but again is out of the scope of this paper.
@@ 4.3 Importance of feature engineering
Feature engineering plays a crucial role in predicting the Eddington tensor accurately. In this work, we incorporate many combined features indeed as listed in Table 1. It is possible because LightGBM is very efficient particularly in training, ignoring irrelevant features automatically. As explained in section 3.2, we find that the flux factor and the non-local are the two most important features for the improvement of accuracy, which we demonstrate here.
In Figure 9, we show how MAE gets better from the one for the model with only the basic features included. The MAE for the model with the basic features alone is actually larger than that for the M1 closure method for most of the neutrino energy. The MAE becomes smaller, however, once the combined features are incorporated, particularly for energies close to the mean energy. It is further improved if the non-local and differential features are taken into account in addition. The same trend is observed for the Eddington tensor $k_{\rm{FP}}^{11}$ itself as a function of radius as demonstrated in Figure 10.
What we learned from these experiments on feature engineering may be three-fold. Firstly, the local low-order moments, the inputs for the ordinary closure methods, are not sufficient to reproduce the Eddington tensor accurately. We find that the information on the background matter such as the baryon density or the optical depth is particularly important. Secondly, the non-local information on the neutrino distribution is probably the most crucial. Although their incorporation in LightGBM is rather straightforward, it will not be so easy for the analytic closure relation to take them into account. Thirdly, finding nice combinations of the basic features is also an important thing. Our LightGBM model is useful in this respect since we can try many possible combinations in a relatively short time. The new features so obtained may be employed in other machine learning models.
Figure 9: The same as the bottom left figure in figure 3 except there are some additional plots: the black one is the LightGBM prediction with basic features (ll. 1–6 of Table 1) and the orange line with combined features.
Figure 10: LightGBM prediction of $k_{\rm{FP}}^{11}$ with different features $(200\,{\rm ms})$. The blue line corresponds to the Boltzmann-radiation-hydrodynamics code and the green line to the M1 closure method. The black line is the LightGBM prediction with basic features (ll. 1–6 of Table 1), the orange line with combined features, and the red line with all engineered features: combined features, non-local features and differential features.
Figure 11: Probability distributions of fluid velocity (”Fluid velocity (r)”) and the collinearity parameter(”collinearity parameter”) for neutrinos with the energy of 1.0 MeV, 8.5 MeV, 18.8 MeV, and 54.0 MeV. Each row corresponds to one of these neutrino energies. The data sets at 300 ms and 200 ms post-bounce are colored blue and orange, respectively. The histograms and KDE (Kernel Density Estimation) plots illustrate the statistical difference between these two data sets for each neutrino energy level. The value of the KL-divergence is given at the plots to quantify the extent of the differences.
5 Conclusion & Discussion
@@ Acknowledgments
We acknowledge Hiroki Nagakura and Keiya Hirashima for fruitful discussions. This work was supported by KAKENHI Grant Numbers JP21K13913, 21H01083. This work was also supported by MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Toward a unified view of the universe: from large scale structures to planets). S. T. is supported by International Graduate Program of Innovation for Intelligent World. S. Y. is supported by Institute for Advanced Theoretical and Experimental Physics, Waseda University and the Waseda University Grant for Special Research Projects (project number: 2023-C141, 2024-C56, 2024-Q014). We acknowledge the high-performance computing resources of the K-computer / the supercomputer Fugaku provided by RIKEN, the FX10 provided by Tokyo University, the FX100 provided by Nagoya University, the Grand Chariot provided by Hokkaido University, and Oakforest-PACS / Wisteria Odyssey provided by JCAHPC through the HPCI System Research Project (Project ID: hp130025, hp230204, hp240219, hp240264, 140211, 150225, 150262, 160071, 160211, 170031, 170230, 170304, 180111, 180179, 180239, 190100, 190160, 200102, 200124, 210050, 220047, 220223, 230056, 230270, 240041) for producing and processing the supervisor data.